# bibliography.bib

### v3

@@ Abstract
1 Introduction
@@
$T(x)\quad\text{for ``$x$ is true''}$”
$F(x)\quad\text{for ``$x$ is false''.}$”.
Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting.
@@
$\mathbf{Ground}$ $\quad(\lozenge T(x)\wedge\lozenge F(x))\to\lozenge N(x)$
@@
Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the normal modal logic that results from adjoining these axioms to $\mathbf{S5}$. The following fine-grained soundness and completeness theorem characterizes the modal theory of $\mathbf{FIX}$:
@@
For all formulas $\varphi$ with $n$ variables, the following are equivalent:
@@
This completeness theorem engenders an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we extract the following answer to our motivating question:
@@
2 Kripke’s Theory of Truth
@@
2.1 Strong Kleene Logic
@@
2.2 Classification of Sentences
Kripke introduces a large collection of 3-valued (true, false, neither) models for $\mathcal{L}_{T}$. These 3-valued models are typically called “fixed points,” since Kripke defines them as the fixed points of a certain function (namely, the Kripke jump). The details of this function are not important for us, but the “fixed points” terminology is ubiquitous, so we will continue to use it.
@@ 2.2 Classification of Sentences
  2. 2.
    $w$ is consistent, i.e., there is no $\varphi$ such that $w\Vdash_{\mathsf{K3}}\varphi$ and $w\Vdash_{\mathsf{K3}}\neg\varphi$.
  3. 3.
@@
For any formula $\varphi(x)$ in the language of $\mathcal{L}_{T}$, we can formulate an $\mathcal{L}_{T}$ sentence $\psi$ which “says” $\varphi(\ulcorner\psi\urcorner)$. In particular, $\psi$ and $\varphi(\ulcorner\psi\urcorner)$ have the same semantic value in each fixed point. For instance, there are $\mathcal{L}_{T}$ sentences $\lambda,\tau,\gamma$ such that:
@@
There are infinitely many logically independent truth-teller sentences. That is, there is an infinite collection $T$ of sentences such that:
  • For all $\tau\in T$, $\tau$ has the same semantic value as $\mathsf{True}(\ulcorner\tau\urcorner)$ in every fixed point.
  • For any sequence $\tau_{1},\dots,\tau_{n}$ of sentences from $T$ and any assignment of true, false, and neither to $\tau_{1},\dots,\tau_{n}$, there is a fixed point realizing exactly that truth-value assignment.
@@ 3 Modal Systems
    $V^{\star}_{1}(x)$ is the set of models in which $x^{\star}$ is true.
@@
    $V^{\star}_{2}(x)$ is the set of models in which $x^{\star}$ is false.
@@
4 Single Variable Definability
@@ 4.2 Counting
$B(x)$
$T(x)\wedge F(x)$
$T_{only}(x)$
$T(x)\wedge\neg F(x)$
$F_{only}(x)$
$\neg T(x)\wedge F(x)$
$N(x)$
$\neg T(x)\wedge\neg F(x)$
$\mathbf{S5}$ has 4 extensional 1-isolators and 15 intensional 1-isolators, namely, all formulas of the form:
@@
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $4\times 8=32$ many 1-isolators for $\mathbf{S5}$. Hence, the number of formulas up to provable equivalence is exactly $2962^{32}=4{,}294{,}967{,}$ 296. ∎
@@
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $3+3+4=10$ many 1-isolators for $\mathbf{S5[Con]}$. Hence, the number of formulas up to provable-equivalence is exactly $10242^{10}=$ 1024. ∎
4.3 Completeness
@@ 4.4 Descriptive Complexity
Every $\mathbf{FIX}$-definable is definable by an intensional 1-formula. This is because if $\varphi$ $\mathbf{FIX}$-defines a set $\mathcal{A}$, then so does $\Box\varphi$. We have already seen that each intensional 1-formula is $\mathbf{S5[Con,Ground]}$-equivalent to a disjunction of intensional 1-isolators. So, pushing quantifiers in, we see that every modal definition is equivalent to a formula of the following form:
@@
4.5 Indiscernibility
The previous result demonstrates the non-definability of a great number of sets, but it is a coarse tool. For instance, descriptive complexity considerations do not yield the non-definability of the arithmetic truths, since they form a $\Delta^{1}_{1}$ set, but the the arithmetic truths are not definable.
We will now prove the indefinability of the intrinsic sentences in another way. Informally, sets $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible just in case every formula that applies to the $\mathcal{A}$s applies to the $\mathcal{B}$s and vice-versa. We will prove that the intrinsic sentences and the inevitable sentences are $\mathbf{FIX}$-indiscernible, whence the set of intrinsic sentences—the smaller of the two sets—cannot be definable.
@@ 4.5 Indiscernibility
Suppose, for the sake of contradiction, that (2) fails. Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$. So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{4}(x)$. Let’s break into cases.
@@ 5.1 Counting Formulas
5.1.1 The Bare Modal System
@@
So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
5.1.2 With the Consistency Axiom
@@
So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
5.2 Matrices
@@
Table 1: All extensional 2-isolators.
@@
Table 2: Standard numbering of the $3\times 3$ matrix.
@@
Table 3: A proper subset of the extensional 2-isolators.
@@
Table 4: A proper subset of the extensional 2-isolators.
@@
Table 5: A proper subset of the extensional 2-isolators.
@@
Table 6: A proper subset of the extensional 2-isolators.
@@ 5.3 The Prime Conditions
$\big{(}\lozenge N(x_{1})\wedge N(x_{1})\wedge\lozenge N(x_{2})\big{)}\to\lozenge\big{(}N(x_{1})\wedge N(x_{1})\wedge N(x_{2})\big{)}$
@@
In the previous subsection we identified intensional $2$-isolators with subsets of the $3\times 3$ matrix. For the $n$-variable case, we must turn from matrices to tensors, which are generalizations of matrices beyond two dimensions. In particular, will be interested in $3^{n}$ tensors, i.e., $\underbrace{3\times\ldots\times 3}\limits_{\mbox{\scriptsize$n$ times}}$ matrices.
Let $[3]^{n}$ be the set of triples from the set $\{1,2,3\}$. This is the set of positions in a $3^{n}$ tensor. We can no longer speak only of rows and columns. Instead, we speak in terms of slices. The rows are the 1-slices, the columns are the 2-slices, and so on. How do we speak of the first row, the second row, the third column, etc.? The 1st row is the 1st layer in the 1st slice, the third column is the 3rd layer in the second slice, and so on. We introduce the following notation.
@@
For $k\leq 3$ and $j\leq n$, the $k^{th}$-layer of the $j^{th}$ slice (written $j_{k}$) is
$\{a\in[3^{n}]\mid\exists b_{1},\dots,\exists b_{n-1}\;a=\langle b_{1},\dots b_{j-1},k,b_{j},\dots,b_{n-1}\rangle\}$
@@
  3. 3.
    For all $l\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
@@ 5.4 The Main Lemma
Suppose that $\varphi\in\mathcal{S}$. Note that $\varphi$ is a conjunction of the form $\varphi_{1}(x_{1})\wedge\dots\wedge\varphi_{n}(x_{n})$. Note also that $\Phi^{-1}(\varphi)$ is some number $k\leq 3^{n}$. Since $C_{k}$ is a Boolean statement about independent truth-tellers, there is some $w$ such that $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$ (see Remark 2.2).
@@
$A^{i}$ is a conjunction of conditionals with antecedent $C_{j}$ such that for $j\neq k$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$. Which is to say that for all conjuncts $A^{i}_{j}$ of $A^{i}$ except for $A^{i}_{k}$, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Thus, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{k}$. Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$, we infer that, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$.
@@
Case 1: Suppose $\mathbb{N},w\uparrow_{\mathsf{K3}}C_{s}$ for all $s\leq 16$. We claim that $A^{1},\dots,A^{n}$ all have value neither in $w$.
@@
Subcase 2.1: Suppose $s>3^{n}$. $A^{i}$ is a conjunction of conditionals of the form $C_{j}\to\theta^{A}_{i}$ such that for $j\neq s$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Accordingly, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$, where $\theta$ is the consequent of $A^{i}_{s}$. Since $s>3^{n}$, $\theta$ is $\lambda$. Hence, $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}$. Hence, $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$ and $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\neg A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}N(x_{i})$.
@@
Subcase 2.2: Suppose that $s\leq 3^{n}$. As in the previous subcase, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}\leftrightarrow\theta$ where $\theta$ is the consequent of $A^{i}_{s}$.
@@
Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
@@
$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta.$

### v4

@@ Abstract.
Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matthew Mandelkern for comments on drafts.
1. Introduction
@@
$T(x)\quad\text{for ``$x$ is true.''}$.”
$F(x)\quad\text{for ``$x$ is false.''}$.”
Kripke’s informal modal definitions of groundedness and paradoxicality are straightforwardly formalizable in this setting. [^2] It is is worth noting that there is other work mixing modality and Kripke’s theory of truth, see [stern2014modality, nicolai2021modal, cook2022outline]. However, none of these address the particular questions addressed here.
@@
$\mathbf{Ground}$ $\quad\big{(}\lozenge T(x)\wedge\lozenge F(x)\big{)}\to\lozenge N(x)$
@@
Let $\mathbf{S5[Con,Ground,Min_{n}]}$ be the modal system that results from adjoining these axioms to $\mathbf{S5}$. The following fine-grained soundness and completeness theorem characterizes the modal theory of $\mathbf{FIX}$:
@@
For all formulas $\varphi$ with at most $n$ variables, the following are equivalent:
@@
This completeness theorem emerges from an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we extract the following answer to our motivating question:
@@
2. Kripke’s Theory of Truth
@@
2.1. Strong Kleene Logic
@@
2.2. Classification of Sentences
Kripke introduces a large collection of 3-valued (true, false, neither) models for $\mathcal{L}_{T}$. These 3-valued models are typically called “fixed points,” since Kripke defines them as the fixed points of a certain function (namely, the Kripke jump). The details of this function are not important for us, but the “fixed points” terminology is ubiquitous, so we will continue to use it. For work on the structure of these fixed points see [cantini1989notes, fitting1986notes].
@@ 2.2. Classification of Sentences
  2. (2)
    $w$ is consistent, i.e., there is no $\varphi$ such that $w\vDash_{\mathsf{K3}}\varphi$ and $w\vDash_{\mathsf{K3}}\neg\varphi$ [^4] Like Kripke, we restrict our attention to consistent fixed points. It is now common to study all fixed points without this restriction.
  3. (3)
@@
For any formula $\varphi(x)$ in the language of $\mathcal{L}_{T}$, a standard self-reference construction yields an $\mathcal{L}_{T}$ sentence $\psi$ which “says” $\varphi(\ulcorner\psi\urcorner)$. In particular, $\psi$ and $\varphi(\ulcorner\psi\urcorner)$ have the same semantic value in each fixed point. For instance, there are $\mathcal{L}_{T}$ sentences $\lambda,\tau,\gamma$ such that:
@@
There are infinitely many logically independent truth-teller sentences. Hence, there is an infinite collection $T$ of sentences such that for any sequence $\tau_{1},\dots,\tau_{n}$ of sentences from $T$ and any assignment of true, false, and neither to $\tau_{1},\dots,\tau_{n}$, there is a fixed point realizing exactly that truth-value assignment.
@@ 3. Modal Systems
    $V^{\star}_{1}(x)$ is the set of worlds in which $x^{\star}$ is true.
@@
    $V^{\star}_{2}(x)$ is the set of worlds in which $x^{\star}$ is false.
@@
4. Single Variable Definability
@@ 4.2. Counting
$B(x)$ $\quad T(x)\wedge F(x)$
$T_{only}(x)$ $\quad T(x)\wedge\neg F(x)$
$F_{only}(x)$$\quad\neg T(x)\wedge F(x)$
$N(x)$$\quad\neg T(x)\wedge\neg F(x)$
$\mathbf{S5}$ has 4 extensional 1-isolators and 15 intensional 1-isolators, namely, all 1-formulas of the form:
@@
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $4\times 8=32$ many 1-isolators for $\mathbf{S5}$. Hence, the number of 1-formulas up to provable equivalence is exactly $2962^{32}=4{,}294{,}967{,}$ 296. ∎
@@
A 1-isolator is a consistent conjunction of an extensional 1-isolator and an intensional 1-isolator. So there are $3+3+4=10$ many 1-isolators for $\mathbf{S5[Con,Ground]}$. Hence, the number of formulas up to provable-equivalence is exactly $10242^{10}=$ 1024. ∎
4.3. Completeness
@@
Before continuing, we note that the normal form theorem also yields an exact characterization of the number of FIX-definable sets.
@@ 4.4. Descriptive Complexity
Every $\mathbf{FIX}$-definable set is definable by an intensional 1-formula. This is because if $\varphi$ $\mathbf{FIX}$-defines a set $\mathcal{A}$, then so does $\Box\varphi$. We have already seen that each intensional 1-formula is $\mathbf{S5[Con,Ground]}$-equivalent to a disjunction of intensional 1-isolators. So, pushing quantifiers in, we see that every modal definition is equivalent to a formula of the following form:
@@
4.5. Indiscernibility
The previous result demonstrates the non-definability of a great number of sets. Nevertheless, it is a coarse tool. We can prove the indefinability of the intrinsic sentences in a more informative way.
Informally, sets $\mathcal{A}$ and $\mathcal{B}$ are $\mathbf{FIX}$-indiscernible just in case every formula that applies to the $\mathcal{A}$s applies to the $\mathcal{B}$s and vice-versa. We will prove that the intrinsic sentences and the inevitable sentences are $\mathbf{FIX}$-indiscernible, whence the set of intrinsic sentences—the smaller of the two sets—cannot be definable.
@@ 4.5. Indiscernibility
Suppose, for the sake of contradiction, that (2) fails. Then there is some $\star$ such that $y^{\star}=\psi$ is inevitable and that $\mathbf{FIX}\nVdash_{\mathbf{V^{\star}}}\varphi(y)$. So $\varphi$ must be missing one of the disjuncts required to define the inevitable sentences, i.e., $\psi_{2}(x)$,…,$\psi_{5}(x)$. Let’s break into cases.
@@ 5.1. Counting Formulas
5.1.1. The Bare Modal System
@@
So there are $4^{n}\times 2^{(4^{n}-1)}$ many $n$-isolators. So the number of $n$-formulas up to provable equivalence is $2^{(4^{n}2^{(4^{n}-1)})}$. ∎
5.1.2. With the Consistency Axiom
@@
So there are $3^{n}\times 2^{(3^{n}-1)}$ many $n$-isolators. So the number of $n$-formulas up to provable equivalence is $2^{(3^{n}2^{(3^{n}-1)})}$. ∎
5.2. Matrices
@@
Table 1. All extensional 2-isolators.
@@
Table 2. Standard numbering of the $3\times 3$ matrix.
@@
Table 3. A proper subset of the extensional 2-isolators.
@@
Table 4. A proper subset of the extensional 2-isolators.
@@
Table 5. A proper subset of the extensional 2-isolators.
@@
Table 6. A proper subset of the extensional 2-isolators.
@@ 5.3. The Prime Conditions
$\big{(}\lozenge N(x_{1})\wedge\lozenge N(x_{1})\wedge\lozenge N(x_{2})\big{)}\to\lozenge\big{(}N(x_{1})\wedge N(x_{1})\wedge N(x_{2})\big{)}$
@@
In the previous subsection we identified intensional $2$-isolators with subsets of the $3\times 3$ matrix. For the $n$-variable case, we must turn from matrices to tensors, which are generalizations of matrices beyond two dimensions. In particular, we will be interested in $3^{n}$ tensors, i.e., $\underbrace{3\times\ldots\times 3}\limits_{\mbox{\scriptsize$n$ times}}$ matrices.
Let $[3]^{n}$ be the set of $n$-tuples from the set $\{1,2,3\}$. This is the set of positions in a $3^{n}$ tensor. We can no longer speak only of rows and columns. Instead, we speak in terms of slices. The rows are the 1-slices, the columns are the 2-slices, and so on. How do we speak of the first row, the second row, the third column, etc.? The 1st row is the 1st layer in the 1st slice, the third column is the 3rd layer in the second slice, and so on. We introduce the following notation.
@@
For $k\leq 3$ and $j\leq n$, the $k^{th}$ layer of the $j^{th}$ slice (written $j_{k}$) is
$\{a\in[3]^{n}\mid\exists b_{1},\dots,\exists b_{n-1}\;a=\langle b_{1},\dots b_{j-1},k,b_{j},\dots,b_{n-1}\rangle\}$
@@
  3. (3)
    For all $t\leq n$, for any $t$ slices $s^{1},\dots,s^{t}$, if $\mathcal{S}$ intersects $s^{i}_{3}$ for each $i\leq t$, then $\mathcal{S}$ intersects $s^{1}_{3}\cap\dots\cap s^{t}_{3}$.
@@ 5.4. The Main Lemma
Suppose that $\varphi\in\mathcal{S}$. Note that $\varphi$ is a conjunction of the form $\varphi_{1}(x_{1})\wedge\dots\wedge\varphi_{n}(x_{n})$. Note also that $\Phi^{-1}(\varphi)$ is some number $k\leq 3^{n}$ (by $\Phi^{-1}$ I mean the inverse of $\Phi$; so $\Phi^{-1}(\varphi)$ is just the number that $\Phi$ maps to $\varphi$). Since $C_{k}$ is a Boolean statement about independent truth-tellers, there is some $w$ such that $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$ (see Remark 2.2).
@@
$A^{i}$ is a conjunction of conditionals with antecedent $C_{j}$ such that for $j\neq k$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$. Which is to say that for all conjuncts $A^{i}_{j}$ of $A^{i}$ except for $A^{i}_{k}$, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Thus, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{k}$. Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{k}$, we infer that $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta$.
@@
Case 1: Suppose $\mathbb{N},w\uparrow_{\mathsf{K3}}C_{s}$ for all $s\leq 2^{k}$. We claim that $A^{1},\dots,A^{n}$ all have value neither in $w$.
@@
Subcase 2.1: Suppose $s>3^{n}$. $A^{i}$ is a conjunction of conditionals of the form $C_{j}\to\theta^{A}_{i}$ such that for $j\neq s$, $\mathbb{N},w\mathrel{\mathchoice{\reflectbox{$\displaystyle\vDash$}}{\reflectbox{$\textstyle\vDash$}}{\reflectbox{$\scriptstyle\vDash$}}{\reflectbox{$\scriptscriptstyle\vDash$}}}_{\mathsf{K3}}C_{j}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{j}$. Accordingly, $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\uparrow_{\mathsf{K3}}\theta$, where $\theta$ is the consequent of $A^{i}_{s}$. Since $s>3^{n}$, $\theta$ is $\lambda$. Hence, $\mathbb{N},w\uparrow_{\mathsf{K3}}A^{i}$. Hence, $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$ and $\mathbb{N},w\nVdash_{\mathsf{K3}}\mathsf{True}(\ulcorner\neg A^{i}\urcorner)$. Since $x_{i}^{\star}=A^{i}$, $\mathbf{FIX},w\Vdash_{V^{\star}}N(x_{i})$.
@@
Subcase 2.2: Suppose that $s\leq 3^{n}$. As in the previous subcase, $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta$ where $\theta$ is the consequent of $A^{i}_{s}$.
@@
Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
@@
$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\Leftrightarrow\mathbb{N},w\vDash_{\mathsf{K3}}\theta.$